Assume that the risk-free rate of return is 4% and the market risk premium (i.e., Rm - Rf ) is 8%. If use the Capital Asset Pricing Model (CAPM) to estimate the expected rate of return on a stock with a beta of 1.28, then this stock’s expected return should be ____.

 

A)

10.53%

 

B)

14.24%

 

C)

23.15%

 

D)

6.59%

What is the beta of a stock with an expected return of 10%, if Treasury bills yield 4% and the market return (i.e., Rm) is 10%?

 

A)

3.0

 

B)

2.0

 

C)

1.0

 

D)

0.2

A company has a project with initial investment is $40,000. It will generate $15,000 annually for the next four years. Assume that this company and its project have a beta of 2.0, the risk-free rate of return (i.e., Rm) is 2%, and the market return (i.e., Rm) is 7%?. How much is the NPV of this project?   [Hint: As discussed, the CAMP model can be used to estimate discount rate (r) in the NPV analysis equation].

 

A)

5,555

 

B)

3,333

 

C)

4,444

 

D)

6,666
Answer #1

(a) Risk-Free Rate = Rf = 4% and Market Risk Premium = MRP = 8 %, Beta = 1.28

Using CAPM, the Expected Stock Return = 4 + 1.28 x 8 = 14.24 %

Hence, the correct option is (C)

(b) Expected Return = 10%, Risk-Free Rate = 4% and Market Return = 10 %, Let the beta be B

Therefore, 10 = 4 + B x (10-4)

B = (10-4) / (10-4) = 1

Hence, the correct option is (C)

(c) Initial Investment = $40000, Annual Payouts = $ 15000, Tenure = 4 years, Risk-Free Rate = 2%, Market Return = 7% and Beta = 2

Using CAPM, Expected Rate of Return = 2 + 2 x (7-2) = 12 %

Therefore, NPV = 15000 x (1/0.12) x [1-{1/(1.12)^(4)}] - 40000 = $ 5560.24 ~ $5555

Hence, the correct option is (A)